Probability

30)One card is drawn from a well shuffled deck of 52 cards.

is equally likely, calculate the probability that the card will be


not a black card.
Answer #1

The event card drawn is not a black card may be denoted as c or not c .we know that P( not c)=1-P(c)=1-1/2=1/2

Therefore, probability of not a black card =1)

2

answered by: Vijayan